# OTAD: An Optimal Transport-Induced Robust Model for Agnostic Adversarial Attack

### v1

@@ Abstract
Index Terms:
Adversarial defense, Lipschitz network, optimal transport, convex integration problem
@@ I Introduction
Various strategies have been proposed to enhance existing models [8]. These strategies include adversarial training [9, 10, 11], where adversarial examples are generated during training and added to the training set. However, these modified models often exhibit vulnerabilities against strong adversaries [12, 13, 14, 15], primarily because DNNs require large gradients to represent their target functions and attacks can always take advantage of the gradients to construct adversarial examples. To stop playing this cat-and-mouse game, a growing body of studies have focused on certified robustness.
@@
To further improve the performance of OTAD, we extend it with various architectures and metric learning. First, we adapt OTAD to Transformer-based architecture such as ViT [28], as Transformers employ residue connections in forward propagation, and the invariant dimension of the feature aligns well with the optimal transport setting. Second, since finding neighbors is a critical step in OTAD, the $l_{2}$ distance may not effectively characterize the similarity of data closing to a manifold in high dimensional space. Consequently, we explore metric learning to find more suitable neighbors and explore the trade-off between accuracy and vulnerability in such cases. We implement various experiments to test the proposed OTAD model and its variants under different settings. Empirical results demonstrate superior performance to adversarial training methods and Lipschitz networks across diverse datasets.
The rest of this paper is organized as follows. In section 2, we present the background of our method: the optimal transport theory and regularity of the optimal transport map. In section 3, we develop the Optimal Transport-based Adversarial Defense model (OTAD) and its implementation details. In section 4, we perform extensive experiments to demonstrate the defense ability of OTAD.
@@ II-A Adversarial training
where $p(x,y)$ is the joint distribution of data and labels, $\mathcal{B}(x)$ is a neighborhood of $x$. However, it is impossible to search the whole region since the loss surface of networks is complex. To approximate it, adversarial training methods [9, 10, 11, 29, 30, 31, 32, 33] add adversarial examples (often found by gradient descent) to the training set during training. After training, the model is robust to the type of attack chosen in the training process. The defense can still be ‘broken’ by stronger adversaries [12, 13, 14, 15]. To escape the cat-and-mouse game, training neural networks with bounded Lipschitz continuity has been considered a promising way to defend against attacks.
@@ II-B Adversarial purification
Adversarial purification can defend unseen attacks because the generative model $G$ is trained independently from adversarial attacks $n_{ad}$ and classifiers $C$. However, we can still construct adversarial examples by taking the gradients of the generative process. Such adversarial noise will not be purified [37]. The effectiveness of adversarial purification methods is highly related to the performance of generative models, such as energy-based models (EBM)[38], generative adversarial networks (GAN) [39, 40] and auto-regressive generative models[41]. The diffusion model is a rising effective generative model. Diffusion-based adversarial methods have achieved competitive performance on large-scale image datasets [42, 43, 37]. For small datasets (e.g., single-cell gene expression data) with specific noise or missing values, it can be hard to train an effective generative model, indicating that adversarial purification may not be a general method for various datasets.
@@ II-C Lipschitz networks and random smoothing
Training neural networks under a Lipschitz constraint puts a bound on how much its input can change in proportion to a change in its input. Existing methods fall into three categories: soft regularization, hard constraints for weights, and specifically designed activations.
Regularizations such as penalizing the Jacobian of the network [16, 17, 18] can constrain the Lipschitz constant locally but don’t provably enforce a global Lipschitz constraint. Thus, adding such regularizations cannot solve the fragility of adversarial attacks on neural networks.
@@ III-A Optimal transport
Optimal transport theory provides valuable tools for quantifying the closeness between probability measures, even when their supports do not overlap. $\mathcal{P}_{2}(\mathbb{R}^{d})$ denotes the set of Borel probability measures with finite second-order moment. For two probability measures $\mu$,$\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, let $\Pi(\mu,\nu)$ denote the set of all joint distributions $\pi(x,y)$ whose marginals are $\mu$ and $\nu$ respectively. The Wasserstein distance is defined as the solution of the Kantorovich problem [45]:
[…]
@@ IV-A ResNet-based OTAD
where $R_{k}$ denotes the shallow network inside the $k$-th block of $R$. Let $R(x)$ be the last-layer features, i.e., $R(x)=x^{m}$. Here $\{x^{k}_{1},x^{k}_{2},\cdots,x^{k}_{n}\}$ can be viewed as particles of the distribution $P_{k}$, then the sequential of distribution $\{P_{k}\}_{k=1}^{m}$ is a discretization of a continuity equation
@@
According to the Benamou-Brenier formula in (6), when $p_{0}$ and $p_{1}$ are fixed, the Wasserstein geodesic curve induced by optimal transport map can be recovered by minimizing the energy
[…]
A discretization of this energy concerning the ResNet $R(\cdot)$ is
@@
Then we obtain the feature $z^{\prime}$ used to classify $x^{\prime}$. If $\{(x,z)|x\in N_{K}(x^{\prime}),z=T(x)\}$ is not $\mathcal{F}_{l,L}$-integrable, we repeat the procedures above with smaller $l$ and larger $L$ until we find feasible values. The solution always exists when $l$ is sufficiently small and $L$ sufficiently large. The training and testing scheme of ResNet-based OTAD is summarized in Algorithms 1 and 2.
1:data $\{x_{i}\}_{i=1}^{n}$, labels $\{y_{i}\}_{i=1}^{n}$, hyperparameter $\lambda$
@@ IV-C Neural network for CIP
Specifically, the inputs of CIP are the embeddings of test data $\tilde{x}$, its neighbors in training set $N_{K}(\tilde{x})$ and the features corresponding to the neighbors $Z(\tilde{x})=\{\tilde{z}_{i}|\tilde{x}_{i}\in N_{K}(\tilde{x})\}$. The output is the solution of the QCP solver, i.e., the estimated feature $\tilde{z}$ of test data. Since the attention block in the Transformer can learn the complex relation between tokens, we train a Transformer to align the inputs and outputs of the CIP, using the solutions from the QCP solver as the training data. We use the MSE loss to fit the QCP solver’s solutions. The resulting Transformer is called CIP-net. Assume we have a training set $S$, here $N_{K}(\tilde{x})$ is the neighborhoods set in $S$ excluded $\tilde{x}$, then we can train the CIP-net by
@@
Let $A^{(r)}=K^{(r)}Q^{(r)\top}/\sqrt{D/R}\in\mathbb{R}^{D\times D}$, consider the map $\tilde{f}$ from $\mathbb{R}^{N\times D}$ to $\mathbb{R}^{N\times D}$, where $f^{(r)}=\tilde{f}^{(r)}V^{(r)}$,
$\tilde{f}:=\operatorname{softmax}\left(XA^{\top}X^{\top}\right)X$
The Jacobian of $\tilde{f}$ is
$J_{\tilde{f}}=\left[\begin{array}[]{ccc}J_{11}&\ldots&J_{1N}\\ \vdots&\ddots&\vdots\\ J_{N1}&\ldots&J_{NN}\end{array}\right]\in\mathbb{R}^{ND\times ND}$
where $J_{ij}=X^{\top}P^{(i)}\left[E_{ji}XA^{\top}+XA\delta_{ij}\right]+P_{ij}I$. Here, $E_{ij}\in\mathbb{R}^{N\times N}$ is a binary matrix with zeros everywhere except at the $(i,j)$-th entry, $\delta_{ij}$ is the Kronecker delta, and $P^{(i)}:=\operatorname{diag}(P_{i:})-P_{i:}^{\top}P_{i:}$. The vector $P_{i:}$ is defined as $\operatorname{softmax}(XA^{\top}x_{i})$.
@@
Hence, $\operatorname{Lip}_{2}(\tilde{f})\leq\dfrac{\left\|KQ^{T}\right\|_{2}}{\sqrt{D/R}}M^{2}(\sqrt{N}+1)+\sqrt{N}$. For single-head dot-product self-attention $f^{(r)}=\tilde{f}^{(r)}V^{(r)}$, any parameter matrix $\|A\|_{2}\leq\|A\|_{F}\leq M_{\theta}$, then
@@
The theorem shows that the local Lipschitz constant of self-attention is bounded by the norm of the parameter matrix. Consequently, training with weight decay can help reduce the Lipschitz constant of self-attention, resulting in a more robust model.
@@ IV-D Finding better neighbors through metric learning
In this paper, we choose ResNet and one attention block in deep metric learning for neighbor searching. Our results demonstrate that deep metric learning can significantly enhance the performance of OTAD in handling complex data. However, DML-net itself remains vulnerable to adversarial attacks. For instance, an untargeted Projected Gradient Descent (PGD) attack can disrupt the learned features, making OTAD worse. In a word, we can always search for a robust DML-net $\psi_{\theta}(\cdot)$ [58] for neighbor searching to improve the performance of OTAD while defending against adversarial attacks. If $\psi_{\theta}(\cdot)$ is not a robust model, we face a trade-off: better neighborhoods but increased vulnerability.
@@ V Experimental results
In this section, we evaluate OTAD and its extensions. For OTAD, we train it on datasets of diverse data types, e.g., image, single-cell transcriptomics, and industrial tabular data. We compare the performance of OTAD against adversarial attacks on these datasets. We change the hyperparameter in OTAD to show its effect on the Lipschitz constant. For OTAD-T, we test it with DML-net on more complex datasets against adversarial attacks and evaluate its robustness. For OTAD-T-NN, we show its inference time and robustness against gradient-free and gradient attacks. We also compare our method with the commonly used $k$ nearest neighborhood-based method (KNN). Since our model is constructed based on the theoretical understanding of residual networks [26], we also test similar models based on plain networks without residual connections for comparison. Finally, we discuss the limitation of our model on synthetic data.
@@ V-A Experimental setup
For OTAD, we set the weight decay hyper-parameter to $5\times 10^{-4}$ since we empirically found that weight decay can help ResNet learn the geodesic curve. We used an SGD optimizer with 0.9 momentum on the image, single-cell transcriptomics, and industrial data. We used the fully connected (FC) ResNet with 5, 4, and 10 ResNet blocks on the three datasets, respectively. The intermediate dimension in each block is set to be the same as the dimension of inputs.
For OTAD-T on CIFAR10 and TinyImageNet, we train Vision Transformers (ViT) with 6 blocks and 8 heads. The input images are divided into 64 patches mapped to a 256-dimensional embedding. The models are optimized using the AdamW optimizer with a cosine annealing learning rate scheduler.
For the metric in neighborhood searching, we primarily use the $l_{2}$ distance. We choose shallow networks as DML-nets to obtain more similar neighbors while maintaining robustness, including 3-block ResNet, 2-layer CNN, 1-block ViT, and an attention block. Specifically, we choose the ResNet-18 pre-trained by ImageNet-1K as DML-net for TinyImageNet. DML-nets are trained/finetuned using the triplet loss, focusing on the difference between the anchor-positive and anchor-negative distances of the learned features.
@@ V-B Evaluation of attacks
We mainly consider the $l_{2}$ attack in this paper as the $l_{2}$ distance is adopted in the construction of OTAD. Let $\epsilon$ be the $l_{2}$ norm of the adversarial perturbations. Since the output of OTAD is obtained by solving the QCP, we cannot take the derivative of the output with respect to the input directly as neural networks. We choose gradient-free attacks to test the models. We introduce three attacks to evaluate our OTAD model on classification tasks:
@@
Backward Pass Differentiable Approximation (BPDA) [60] composited with PGD attack [30] BPDA is typically used when there exists an optimization loop or non-differentiable operations in defense modules. OTAD involves an optimization loop for solving the convex integration problem. Thus, the back-propagation of this part is approximated by the original network. We run 100 iterations of the PGD attack.

### v2

@@ I Introduction
Various strategies have been proposed to enhance the robustness of existing models [8]. These strategies include adversarial training [9, 10, 11], where adversarial examples are generated during training and added to the training set. However, these modified models often exhibit vulnerabilities against strong adversaries [12, 13, 14, 15], primarily because DNNs require large gradients to represent their target functions and attacks can always take advantage of the gradients to construct adversarial examples. To stop playing this cat-and-mouse game, a growing body of studies have focused on certified robustness.
@@
To further improve the performance of OTAD, we extend it with various architectures and metric learning. As Transformers employ residue connections in forward propagation, and the invariant dimension of the feature aligns well with the optimal transport setting, we adapt OTAD to Transformer-based architecture such as ViT [28]. Finding neighbors is a critical step in OTAD, and the $l_{2}$ distance may not effectively characterize the similarity of data closing to a manifold in high dimensional space. Consequently, a more general version of OTAD finds neighbors with learnable metrics. We explore metric learning to find more suitable neighbors and explore the trade-off between accuracy and vulnerability in such cases. Besides, we can randomly choose a training subset for the neighborhood search process to reduce the memory and computational cost. Thus, OTAD can be scaled to large complex datasets like ImageNet. We implement various experiments to test the proposed OTAD model and its variants under different settings. Empirical results demonstrate superior performance to adversarial training methods and Lipschitz networks across diverse datasets.
The rest of this paper is organized as follows. In section 2, we introduce some related works about adversarial defense. In section 3, we present the background of our method: the optimal transport theory and regularity of the optimal transport map. In section 4, we develop the Optimal Transport-based Adversarial Defense model (OTAD) and its implementation details. In section 5, we perform extensive experiments to demonstrate the defense ability of OTAD.
@@ II-A Adversarial training
where $p(x,y)$ is the joint distribution of data and labels, $\mathcal{B}(x)$ is a neighborhood of $x$. However, it is impossible to search the whole region since the loss surface of networks is complex. To approximate it, adversarial training methods [9, 10, 11, 29, 30, 31, 32, 33, 34, 35] add adversarial examples (often found by gradient descent) to the training set during training. After training, the model is robust to the type of attack chosen in the training process. The defense can still be ‘broken’ by stronger adversaries [12, 13, 14, 15]. To escape the cat-and-mouse game, training neural networks with bounded Lipschitz constant has been considered a promising way to defend against attacks.
@@ II-B Adversarial purification
Adversarial purification can defend unseen attacks because the generative model $G$ is trained independently from adversarial attacks $n_{ad}$ and classifiers $C$. However, we can still construct adversarial examples by taking the gradients of the generative process. Such adversarial noise will not be purified [39]. The effectiveness of adversarial purification methods is highly related to the performance of generative models, such as energy-based models (EBM)[40], generative adversarial networks (GAN) [41, 42] and auto-regressive generative models[43]. The diffusion model is a rising effective generative model. Diffusion-based adversarial methods have achieved competitive performance on large-scale image datasets [44, 45]. For small datasets (e.g., single-cell gene expression data) with specific noise or missing values, it can be hard to train an effective generative model, indicating that adversarial purification may not be a general method for various datasets.
@@ II-C Lipschitz networks and random smoothing
Training neural networks under a Lipschitz constraint puts a bound on how much its output can change in proportion to a change in its input. Existing methods fall into three categories: soft regularization, hard constraints for weights, and specifically designed activations.
Regularizations such as penalizing the Jacobian of the network [16, 17, 18] can constrain the Lipschitz condition along some directions but does not provably enforce a local Lipschitz constraint on a $\epsilon$-ball of a data point. Thus, adding such regularizations cannot solve the fragility of adversarial attacks on neural networks.
@@ III-A Optimal transport
Optimal transport theory provides valuable tools for quantifying the closeness between probability measures, even when their supports do not overlap. $\mathcal{P}_{2}(\mathbb{R}^{d})$ denotes the set of Borel probability measures with finite second-order moment. For two probability measures $\mu$$\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, let $\Pi(\mu,\nu)$ denote the set of all joint distributions $\pi(x,y)$ whose marginals are $\mu$ and $\nu$ respectively. The Wasserstein distance is defined as the solution of the Kantorovich problem [47]:
$W_{2}(\mu,\nu):=\left(\inf_{\pi\in\Pi(\mu,\nu)}\int\|x-y\|_{2}^{2}d\pi(x,y)\right)^{1/2}$ (2)
@@ IV-A ResNet-based OTAD
where $R_{k}$ denotes the shallow network inside the $k$-th block of $R$. Let $R(x)$ be the last-layer features, i.e., $R(x)=x^{m}$.
The forward Euler discretization of an ODE
$\frac{dx}{dt}=v(x,t)$
is
$x^{k+1}=x^{k}+\Delta t\cdot v(x^{k},t),$
which is similar to the forward propagation formula of ResNet, omitting the time step $\Delta t$. Thus, ResNet can be viewed as a discretization of an ODE. Furthermore, given data points $\{x_{1},x_{2},\cdots,x_{n}\}$, each point has a measure of $\frac{1}{n}$, the ResNet outputs the same number of points as input. This property is measure preserving. An ODE holding measure preserving property satisfies the continuity equation
@@
where $p_{t}$ is the distribution of $x$ at time $t$ and $v_{t}(x)=v(x,t)$.
If the initial distribution $p_{0}$ and the terminal distribution $p_{1}$ are fixed, infinite ODEs satisfying the boundary condition and continuity equation. The Benamou-Brenier formula said, that the Wasserstein geodesic curve induced by optimal transport map can be recovered by minimizing the energy
$\int_{0}^{1}\|v_{t}\|_{L^{2}(p_{t})}dt$ (12)
As discussed above, ResNet can be viewed as a discretization of the ODE satisfying continuity equation. The energy above also has a discrete version to minimize for ResNet, which is
@@
Then we obtain the feature $z^{\prime}$ used to classify $x^{\prime}$. If $\{(x,z)|x\in N_{K}(x^{\prime}),z=T(x)\}$ is not $\mathcal{F}_{l,L}$-integrable, we repeat the procedures above with smaller $l$ and larger $L$ until we find feasible values. Since the smoothness of Brenier potential cannot be guaranteed, the constant $L$ could be very large. To avoid $L$ larger than tolerance, we detect which constraint is violated in inequalities (15) for current $L$ and delete or substitute the $x_{j}$ corresponding to that inequality. By doing this, we can control the $L$ within tolerance. The training and testing scheme of ResNet-based OTAD is summarized in Algorithms 1 and 2. We use the optimization tool MOSEK [52] to find $\{u_{i}\}_{i=1}^{K}$ satisfying (15) and $z^{\prime}$ by solving (16).
The Simplex method for linear programming in (15) is NP-hard, but since the simplex method works well in most cases and the number of variables in (15) is $K$ (usually set $K=5$ or $K=10$), it doesn’t consume too much time. Solving the convex QCP in (16) consumes most of the time. The complexity of QCP in (16) is $O((K+d)^{3})$, where $K$ is the number of neighbors, and $d$ is the dimension of latent space.
1:data $\{x_{i}\}_{i=1}^{n}$, labels $\{y_{i}\}_{i=1}^{n}$, hyperparameter $\alpha$
@@ IV-C Neural network for CIP
[…]
@@
Let $A^{(r)}=K^{(r)}Q^{(r)\top}/\sqrt{D/R}\in\mathbb{R}^{D\times D}$, consider the map $\tilde{f}$ from $\mathbb{R}^{N\times D}$ to $\mathbb{R}^{N\times D}$, where $f^{(r)}=\tilde{f}^{(r)}V^{(r)}$, ~f:=softmax(XA^⊤X^⊤) X The Jacobian of $\tilde{f}$ is J_~f=[J_11…J_1 N⋮⋱⋮J_N 1…J_N N] ∈R^N D ×N D where $J_{ij}=X^{\top}P^{(i)}\left[E_{ji}XA^{\top}+XA\delta_{ij}\right]+P_{ij}I$. Here, $E_{ij}\in\mathbb{R}^{N\times N}$ is a binary matrix with zeros everywhere except at the $(i,j)$-th entry, $\delta_{ij}$ is the Kronecker delta, and $P^{(i)}:=\operatorname{diag}(P_{i:})-P_{i:}^{\top}P_{i:}$. The vector $P_{i:}$ is defined as $\operatorname{softmax}(XA^{\top}x_{i})$.
@@
Hence, $\operatorname{Lip}_{2}(\tilde{f})\leq\dfrac{\left\|KQ^{T}\right\|_{2}}{\sqrt{D/H}}M^{2}(\sqrt{N}+1)+\sqrt{N}$. For single-head dot-product self-attention $f^{(r)}=\tilde{f}^{(r)}V^{(r)}$, any parameter matrix $\|A\|_{2}\leq\|A\|_{F}\leq M_{\theta}$, then
@@
The theorem shows that the local Lipschitz constant of self-attention is bounded by the norm of the parameter matrix. Consequently, training with weight decay can help reduce the Lipschitz constant of self-attention, resulting in a more robust model. In our setting, the CIP-net is trained on solutions from the QCP solver. Recall the QCP problem (16), whose constraints restrict the feasible region to a small space. This makes predicting the QCP solution with CIP-net more manageable. During training, the prediction error of CIP-net quickly decreases, and the weight decay term is effectively optimized, leading to a CIP-net with a smaller Lipschitz constant and enhanced robustness in OTAD-T-NN.
@@ IV-D Finding better neighbors through metric learning
In this paper, we choose shallow networks as 3-block ResNet or one attention block, and deep networks as ViT-B/16 in deep metric learning for neighbor searching. Our results demonstrate that deep metric learning can significantly enhance the performance of OTAD in handling complex data. However, DML-net itself remains vulnerable to adversarial attacks. For instance, an untargeted Projected Gradient Descent (PGD) attack can disrupt the learned features, making OTAD worse. In a word, we can always search for a robust DML-net $\psi_{\theta}(\cdot)$ [60] for neighbor searching to improve the performance of OTAD while defending against adversarial attacks. If $\psi_{\theta}(\cdot)$ is not a robust model, we face a trade-off: better neighborhoods but increased vulnerability.
@@ V Experimental results
In this section, we evaluate OTAD and its extensions. For OTAD, we train it on datasets of diverse data types, e.g., image, single-cell transcriptomics, and industrial tabular data. We compare the performance of OTAD against adversarial attacks on these datasets. We change the hyperparameter in OTAD to show its effect on the Lipschitz constant. For OTAD-T, we test it with DML-net on more complex datasets against adversarial attacks and evaluate its robustness. For OTAD-T-NN, we show its inference time and robustness against gradient-free and gradient attacks. We also compare our method with the commonly used $k$ nearest neighborhood-based method (KNN). Since our model is constructed based on the theoretical understanding of residual networks [26], we also test similar models based on plain networks without residual connections for comparison. Finally, we discuss the limitation of our model on synthetic data. We have included the results under various perturbation norms as well as the computational overhead of different OTAD variants in the Appendix.
@@ V-A Experimental setup
We evaluate several OTAD variants across different backbones and solver implementations. OTAD refers to the base model built upon a ResNet backbone, while OTAD-T adopts a Transformer backbone. OTAD-T-NN denotes OTAD-T accelerated with a neural-network-based solver for the CIP problem. For neighborhood search in both OTAD-T and OTAD-T-NN, we use either a small ResNet (OTAD-T with ResNet / OTAD-T-NN with ResNet) or an attention block (OTAD-T with attention / OTAD-T-NN with attention) as the DML-net.
We consider three data types: image (MNIST [61], CIFAR10 [62], and ImageNet [63]), single-cell transcriptomics (MERFISH retina [64]), and industrial data (red wine quality [65]). For OTAD, we set the weight decay hyper-parameter to $5\times 10^{-4}$ since we empirically found that weight decay can help ResNet learn the geodesic curve. We use an SGD optimizer with 0.9 momentum on the image (MNIST), single-cell transcriptomics, and industrial data. We use the fully connected (FC) ResNet with 5, 4, and 10 ResNet blocks on the three datasets, respectively. The intermediate dimension in each block is set to be the same as the dimension of inputs.
For OTAD-T on CIFAR10, we train a Vision Transformer (ViT) from scratch with 7 blocks and 12 heads. The input images are divided into 64 patches mapped to a 384-dimensional embedding. The model is optimized using the AdamW optimizer with a cosine annealing learning rate scheduler. On ImageNet, we choose the commonly used ViT architecture: DeiT-S [66]. These models (ResNet or ViT) achieve 98.26%, 90.29%, and 79.2% test accuracy on MNIST, CIFAR10, and ImageNet, respectively.
For the metric in neighborhood searching, we primarily use the $l_{2}$ distance. We choose shallow networks as DML-nets to obtain more similar neighbors while maintaining robustness, including 3-block ResNet, 2-layer CNN, 1-block ViT, and an attention block. Specifically, we choose the ViT-B/16 [28] pre-trained on ImageNet as DML-net for ImageNet. The final classification layer is replaced with a linear projection followed by $L_{2}$ normalization, ensuring the learned features are in a unit sphere for better distance-based comparisons. DML-nets are trained/finetuned using the triplet loss, focusing on the difference between the anchor-positive and anchor-negative distances of the learned features.
@@ V-B Evaluation of attacks
We mainly consider the $l_{2}$ attack in this paper as the $l_{2}$ distance is adopted in the construction of OTAD. Let $\epsilon$ be the $l_{2}$ norm of the adversarial perturbations. We also report the results under various perturbation norms in Appendix A for completeness. Since the output of OTAD is obtained by solving the QCP, we cannot take the derivative of the output with respect to the input directly as neural networks. We choose gradient-free attacks to test the models. We introduce three attacks to evaluate our OTAD model on classification tasks:
@@
Backward Pass Differentiable Approximation (BPDA) [68] composited with PGD attack [30] BPDA is typically used when there exists an optimization loop or non-differentiable operations in defense modules. OTAD involves an optimization loop for solving the convex integration problem. Thus, the back-propagation of this part is approximated by the original network. We mainly compare OTAD with adversarial purification methods against BPDA +PGD. The number of iterations of the PGD attack is 100 on MNIST and 20 on the other datasets. For the non-purification methods, BPDA + PGD attack reduces to standard PGD attack.
@@
For the differentiable OTAD-T-NN variant, we use one of the strongest adversarial attacks, AutoAttack [70], with the default hyperparameters of standard version for evaluation, including three white-box attacks (APGD-CE [70], APGD-DLR [70], and FAB [71]) and a black-box attack (Square Attack [69]).